Consider the window (of area 0.71 m2) of an airplane. If the pressure inside the plane is atmospheric pressure, while the pressure outside is 17 % of Patm, find the magnitude of the force on the window. Ignore the velocity of the airplane (i.e., ignore Bernoulli’s effect).
Answer #1

The pressure outside of the window is, 17%P,tm 17 100 Put = 0.17(1 atm ) 0.17 atm The difference in the inside and outside pr
Answer #2

SOLUTION :


Inside pressure = 1 atmosphere = 101325 Pa


Outside pressure = 0.17 atmosphere = 0.17 * 101325 = 17225.25 Pa 


These pressures act in opposite directions.


So, net pressure on the window = 101325 - 17225.25 = 84099.75 Pa = 84099.75 N/m^2


So, 


Force on the window 

= Net pressure  in N/m^2  * Area in m^2

= 84099.75 * 0.71  N

= 59710.82 N  = 59711 N approx. (ANSWER).

answered by: Tulsiram Garg